OpenStudy (anonymous):

which is equivalent to 3 over squareroot 5?

OpenStudy (anonymous):

Answer choices:\[3 squareroot 5 \over 5\] squareroot15 over 5 square 15 of square 5 3

OpenStudy (anonymous):

\[\frac{ 3 }{ \sqrt{5} }= \frac{ 3\sqrt{5}}{ 5 }\] If you multiply both the numerator and denominator of the first fraction by √5, you will get the second fraction.

OpenStudy (anonymous):

ty